HETAV PATEL

Grade 11,

A body of mass M moving with velocity V makes a head on collision with another body of mass 2m which is initially at rest the loss of kinetic energy of colliding body which has mass m

A body of mass M moving with velocity V makes a head on collision with another body of mass 2m which is initially at rest the loss of kinetic energy of colliding body which has mass m

Grade:11

1 Answers

Arun
5 years ago
let v is the velocity of mass m & v1 is the final velocity of mass 2m &m 
mv = 3mv1 
v1=v/3 loss of energy = (1/2mv^2 - 1/2mv/2^2)/1/2mv^2 
= 8/9